# Boundless Byte Pair Encoding: Breaking the Pre-tokenization Barrier

### v1

@@ 5 Related work
See Appendix D for additional related work.
@@ Appendix A Pre-tokenizer regular expression
Pre-tokenization applies a regular expression (regex) to each document to form pretokens, which are then each tokenized separately. The regex in this section make extensive use of Unicode categories. These are not supported by the re library used by default in python, instead requiring the more powerful regex library. [^10] https://pypi.org/project/regex/ In the notation of regex, Unicode can be divided into the categories given in Table 1.
@@
# B-8, punctuation and symbols
@@
Comparison of pre-tokenization regular expressions (regex)
@@ A.1 Separation of Unicode Classes
One open question with the pre-tokenization regex is how much the Unicode character classes should be kept separate. The GPT2_PATTERN pattern kept them very separate. The GPT4_PATTERN pattern moved away from that with pattern in line F-2 that allowed any single character besides a letter, digit, or line ending to come before a word. It also combined line endings with other characters in F-4. To study if this was a good idea, the Punct pattern (Dagan et al., 2024) returned to only allowing a space before a word with P-1. They found an improvement for PUNCT_PATTERN over GPT4_PATTERN at a vocabulary size of 32k, but no significant difference at a vocabulary size of 80k. Wegmann et al. (2025) also had mixed results on this question. They found the GPT2_PATTERN with the highest separation was best for tasks requiring robustness to language variation, while more mixing was beneficial to tasks requiring sensitivity to language variation.
@@ A.4 Whitespace
[…]
The F-6 pattern matches the first three of the four spaces between Hello and world. Then F-2 picks off the third space in ’ world’. The F-5 matches multiple lines of whitespace ending in a line-ending, and then finally F-6 picks up the remaining trailing whitespace on the final line.
@@ A.6 Contractions
The T-1 and F-1 patterns match English specific contraction endings like 've or 'll. Disliking the English-specific nature, this was omitted from Punct. Pattern B-1 takes a different language-independent approach by keeping the ending of the contraction together with the word as a single pretoken. Any word containing a straight or curly apostrophe internal to the word is matched, like C'est or J'ai.
Appendix B Efficient implementation
In the representation shown in Section 3.2, a document is divided into pretokens, each of which contains one or more tokens. This is simple conceptually and can be used directly for training, but is extremely slow. Tokenizer training routines commonly use the trick of aggregating the pretokens produced over a large training corpus, and use the aggregate counts when calculating pairwise merge counts. Thus the pretoken [ the’] is only ever tokenized once even though it might appear hundreds of thousands of times. This speedup is crucial for performance reasons.  [^12] We have our own implementation of BPE with deletions using this speedup, based on Andrej Karpathy’s minbpe library. https://github.com/karpathy/minbpe
@@ Appendix B Efficient implementation
The second superword representation can only consider a pairwise merge when two adjacent words are unlocked. After a new pretoken is unlocked, the counts of the superword must be updated accordingly. We track the pairwise and single counts of each representation. For speed, we dynamically compute the changes in counts that result from performing each merge. Despite these optimizations, it still took 4.7 CPU days to train our 1GB training dataset due to the lack of aggregation for supermerges.
Appendix C Token distribution at additional vocabulary sizes
Figure 11 shows the same plot as Figure 6 at the vocabulary sizes of 65,536 and 98,304. The trends are largely the same as in Figure 6, with BoundlessBPE having higher counts at the low frequency end of the distribution compared to baselines, and lower frequencies for the most common tokens. Both these are desirable to have a more uniform distribution.
Appendix D Additional related work
@@
Byte level subword tokenization has become a fundamental component in modern NLP, balancing vocabulary size with morphology and out-of-vocabulary handling. Byte Pair Encoding (BPE; Sennrich et al., 2016) iteratively merges the pair of adjacent tokens with the highest count to build a vocabulary. WordPiece (Schuster & Nakajima, 2012) is similar to BPE, except that merges are selected according to their Pointwise Mutual Information (PMI). Kudo (2018) introduced the UnigramLM tokenizer, a top-down approach that starts with a large vocabulary and prunes tokens based on their contribution to sequence likelihood according to a unigram language model.
@@
The pairwise nature of BPE merges mean some tokens added to the vocabulary simply serve as a bridge to a more popular token, but are not often used on their own. Bostrom & Durrett (2020) observed a dead zone of such tokens in both UnigramLM and BPE vocabularies. The proportion of such tokens were found to be higher in the case of BPE, motivating research into their removal and vocabulary refinement. Lian et al. (2024) and Chizhov et al. (2024) proposed strategies to eliminate these low-frequency tokens formed as intermediate steps. We find Chizhov et al. (2024)’s approach particularly compelling due to its direct integration into BPE training. In contrast, Bauwens & Delobelle (2024) focused on a post-processing step to remove merge rules decreasing morphological alignment. We opted against a purely post-processing approach like Cognetta et al. (2024) to maintain tighter control over vocabulary construction during training.
Appendix E Tokenizer overfitting analysis
Figure 12 provides an analysis of the Rényi efficiency of the trained tokenizers over the training and evaluation corpus. In contrast to the results observed on the evaluation corpus Figure 9), the baseline tokenizers exhibit significantly higher Rényi efficiency on the training data. For instance, the Rényi efficiency of the BPE tokenizer reduces by an average of 9% over all vocabulary sizes from being computed on the train to evaluation corpora.
Higher Rényi efficiency on the training set could imply that these tokenizers have learned to represent the training data very effectively, potentially at the expense of generalization to unseen data, suggesting overfitting to some degree. BoundlessBPE, while achieving slightly lower Rényi efficiency on the training data compared to the baselines, demonstrates a more robust performance on the evaluation set, suggesting a better balance between fitting the training data and generalizing to new data.
Figure 12: Rényi efficiency ($\alpha=2.5$) on the training data (left) and evaluation corpus (right). Tokenizers with higher efficiency are generally desirable.

### v2

@@ 6 Related work
See Appendix˜F for additional related work.
@@ Appendix A Pre-tokenizer regular expression
Pre-tokenization applies a regular expression (regex) to each document to form pretokens, which are then each tokenized separately. The regex in this section make extensive use of Unicode categories. These are not supported by the re library used by default in python, instead requiring the more powerful regex library. [^16] https://pypi.org/project/regex/ In the notation of regex, Unicode can be divided into the categories given in Table˜2.
@@
GPT4O_PATTERN = "|".join([
[…]
(?i:’s|’t|’re|’ve|’m|’ll|’d)?", # O-1 word with some lowercase
[…]
(?i:’s|’t|’re|’ve|’m|’ll|’d)?", # O-2 word with some uppercase
[…]
@@
{listing}
Comparison of existing pre-tokenization regular expressions (regex)
@@
# B-8, optional space, punctuation and symbols
@@
BoundlessBPE regular expression (regex)
@@ A.1 Separation of Unicode Classes
One open question with the pre-tokenization regex is how much the Unicode character classes should be kept separate. The GPT2_PATTERN pattern kept them very separate. The GPT4_PATTERN pattern moved away from that with pattern in line F-2 that allowed any single character besides a letter, digit, or line ending to come before a word. It also combined line endings with other characters in F-4. The GPT4O_PATTERN is similar, further breaking words based on case and keeping contractions together. To study if separate Unicode character classes are good idea, the Punct pattern (Dagan et al., 2024) returned to only allowing a space before a word with P-1. They found an improvement for PUNCT_PATTERN over GPT4_PATTERN at a vocabulary size of 32k, but no significant difference at a vocabulary size of 80k. Wegmann et al. (2025) also had mixed results on this question. They found the GPT2_PATTERN with the highest separation was best for tasks requiring robustness to language variation, while more mixing was beneficial to tasks requiring sensitivity to language variation.
@@ A.4 Whitespace
With the example ’Hello world $n \n$ and GPT4_PATTERN we get:
[’Hello’, , world’, $n \n$’, ].
The F-6 pattern matches the first three of the four spaces between Hello and world. Then F-2 picks off the third space in ’ world’. The F-5 matches multiple lines of whitespace ending in a line-ending, and then finally F-6 picks up the remaining trailing whitespace.
@@ A.6 Contractions
The T-1 and F-1 patterns match English specific contraction endings like 've or 'll. Disliking the English-specific nature, this was omitted from Punct. Pattern B-1 is more language-independent. It keeps the ending of the contraction together with the word as a single pretoken. Thus, any word containing a straight or curly apostrophe internal to the word is matched, like C'est or J'ai. Patterns O-1 and O-2 keep the contraction ending with the word.
Appendix B Token distribution at additional vocabulary sizes
Figure˜12 shows the same plot as Figure˜6 at the vocabulary sizes of 65,536 and 98,304. The trends are largely the same as in Figure˜6, with BoundlessBPE having higher counts at the low frequency end of the distribution compared to baselines, and lower frequencies for the most common tokens. Both these are desirable to have a more uniform distribution.
Appendix C Efficient implementation
In the representation shown in Section˜3.2, a document is divided into pretokens, each of which contains one or more tokens. This is simple conceptually and can be used directly for training, but is extremely slow. Tokenizer training routines commonly use the trick of aggregating the pretokens produced over a large training corpus, and use the aggregate counts when calculating pairwise merge counts. Thus the pretoken [ the’] is only ever tokenized once even though it might appear hundreds of thousands of times. This speedup is crucial for performance reasons.  [^19] We have our own implementation of BPE with deletions using this speedup, based on Andrej Karpathy’s minbpe library. https://github.com/karpathy/minbpe
@@ Appendix C Efficient implementation
The second superword representation can only consider a pairwise merge when two adjacent words are unlocked. After a new pretoken is unlocked, the counts of the superword must be updated accordingly. We track the pairwise and single counts of each representation. For speed, we dynamically compute the changes in counts that result from performing each merge. Despite these optimizations, it still took 4.7 CPU days to train our 1GB training dataset due to the lack of aggregation for supermerges. For comparison, training Hugginface BPE on the same data, vocabulary size, and machine takes 59s.
Appendix D Metrics for token distribution uniformity
Zouhar et al. (2023) suggest the use of Rényi efficiency as a metric to measure token distribution uniformity. The equation contains a hyperparameter $\alpha\geq 1$ which controls the exponent on each proportion. They suggest the use of $\alpha=2.5$, which is in the range that strongly penalizes very common tokens. Figure˜13 show the relative Rényi efficiency for various values of $\alpha$. The improvement of BoundlessBPE (4o) over the baselines increases at lower values of $\alpha$. BoundlessBPE (B) outperforms the baselines at lower values of $\alpha$.
Figure 13: Rényi efficiency at various values of $\alpha$
Appendix E Ablation results at additional vocabulary sizes
Table˜3 shows the ablation results at the vocabulary sizes of 65,536 and 98,304.
Vocab Size Regex Pattern PickyBPE Bytes Per Token Rényi Efficiency
65,536 Boundless none 4.488 0.4386
65,536 Boundless original 4.486 0.4386
65,536 Boundless ours 4.480 0.4387
65,536 GPT4o none 4.673 0.4652
65,536 GPT4o original 4.671 0.4653
65,536 GPT4o ours 4.664 0.4653
98,304 Boundless none 4.660 0.4181
98,304 Boundless original 4.658 0.4182
98,304 Boundless ours 4.651 0.4183
98,304 GPT4o none 4.871 0.4439
98,304 GPT4o original 4.869 0.4439
98,304 GPT4o ours 4.861 0.444
Table 3: Effect of regular expression pattern and PickyBPE style on intrinsic measures for vocabulary sizes of 65,536 and 98,304
Appendix F Additional related work
@@
Byte level subword tokenization has become a fundamental component in modern NLP, balancing vocabulary size with morphology and out-of-vocabulary handling. Byte Pair Encoding (BPE; Sennrich et al., 2016) iteratively merges the pair of adjacent tokens with the highest count to build a vocabulary. WordPiece (Schuster & Nakajima, 2012) is similar to BPE, except merges are selected according to their Pointwise Mutual Information (PMI). Kudo (2018) introduced the UnigramLM tokenizer, a top-down approach that starts with a large vocabulary and prunes tokens based on their contribution to sequence likelihood according to a unigram language model.
@@
The pairwise nature of BPE merges mean some scaffold tokens added to the vocabulary simply serve as a bridge to a more popular token, but are not often used on their own. Bostrom & Durrett (2020) observed a dead zone of such tokens in both UnigramLM and BPE vocabularies. The proportion of such tokens were found to be higher in the case of BPE, motivating research into their removal and vocabulary refinement. In Lian et al. (2024) scaffold tokens are marked during a training process, and split into components as a postprocessing step during inference. Chizhov et al. (2024) integrate the deletion step into the BPE training process, where it can affect later merge decisions. Thus an ordered combination of merges and deletions occur during inference. We find Chizhov et al. (2024)’s approach particularly compelling due to its direct integration into BPE training. In contrast, Bauwens & Delobelle (2024) focused on a post-processing step to remove merge rules decreasing morphological alignment. We opted against a purely post-processing approach like Cognetta et al. (2024a) to maintain tighter control over vocabulary construction during training.